James is five years old now and Emma is four years older.
Form a relevant equation and use it to find out how many years it will take until James' and Emma's ages in years, multiplied together, make 725 years.

Answer by nabla(475)

J=5+x
E=9+x
(5+x)(9+x)=725
x^2+14x+45=725
x^2+14x-680=0
(-20 + x) (34 + x)=0
So x=20 or x=-34. x=-34 does not make sense so we exclude it from the answer set.
If x=20, J will be 25, E will be 29.
25*29=725
(The answer was after 20 years from the current time.)